OpenStudy (anonymous):

given the function g(x)=x^2/(x^2+1) find the following: a) State the intervals of increase and decrease in g. b) State the intervals of upward and downward concavity in g. Show the calculations that lead to your conclusions.

myininaya (myininaya):

for a you need to find the critical numbers

myininaya (myininaya):

first we need the derivative to do that

myininaya (myininaya):

\[f'(x)=\frac{(x^2)'(x^2+1)-x^2(x^2+1)'}{(x^2+1)^2}=\frac{2x(x^2+1)-x^2(2x+0)}{(x^2+1)^2}\]

myininaya (myininaya):

\[f'(x)=\frac{2x(x^2+1)-2x^3}{(x^2+1)^2}=\frac{2x(x^2+1-x^2)}{(x^2+1)^2}=\frac{2x}{(x^2+1)^2}\]

myininaya (myininaya):

is that what you got?

myininaya (myininaya):

for f'?

myininaya (myininaya):

now x^2+1 is never 0 so we don't have to worry with f' being undefined but 2x=0 when x=0 so x=0 is a critical number

OpenStudy (anonymous):

yes that is what i got so far

myininaya (myininaya):

ok great

OpenStudy (anonymous):

But then from there i take the derivative of the last equation you got right?

myininaya (myininaya):

so test the areas around x=0 with f' to see if the function is decreasing there or increasing

myininaya (myininaya):

lets do b after we have completely done a

myininaya (myininaya):

unless you totally understand a

OpenStudy (anonymous):

so then lets plug in like -1 & 1 in order to test it?

myininaya (myininaya):

ok those are good numbers

OpenStudy (anonymous):

so we plug in -1 &1 into what equation though?

myininaya (myininaya):

its obvious that f will be decreasing before x=0 since f'<0 and f will be increasing after x=0 since f'>0

myininaya (myininaya):

f'

myininaya (myininaya):

f' tells if a function is increasing or decreasing

myininaya (myininaya):

so you use f' to test for that

OpenStudy (anonymous):

oh okay now i get it! thanks so now we work on b?

myininaya (myininaya):

ok by the way i called your function f instead of g did you notice that?

myininaya (myininaya):

just replace my f's with g's okay?

OpenStudy (anonymous):

yes i did notice that.

OpenStudy (anonymous):

no worries i will fix that.

OpenStudy (anonymous):

yes we do.

myininaya (myininaya):

so anyways we found f' which tells us increasing and decreasing now we need find f'' which tells about f's concavity

myininaya (myininaya):

so \[f''(x)=\frac{(2x)'(x^2+1)^2-2x[(x^2+1)^2]'}{((x^2+1)^2)^2}\]

myininaya (myininaya):

\[f''(x)=\frac{2(x^2+1)^2-2x \cdot 2(x^2+1)(2x)}{(x^2+1)^4}\]

myininaya (myininaya):

is that okay so far?

myininaya (myininaya):

do you understand what i did?

OpenStudy (anonymous):

yes i understand what you did so far

myininaya (myininaya):

\[f''(x)=\frac{2(x^2+1)[(x^2+1)-2x (2x)]}{(x^2+1)^4}=\frac{2(x^2+1)[(x^2+1)-4x^2]}{(x^2+1)^4}\]

myininaya (myininaya):

let me know if you see a mistake

myininaya (myininaya):

\[f''(x)=\frac{2(x^2+1)(-3x^2+1)}{(x^2+1)^4}\]

myininaya (myininaya):

now f'' is 0 when -3x^2+1=0 because x^2+1 is not zero for any real value x

myininaya (myininaya):

\[3x^2=1 => x^2=\frac{1}{3} => x=\pm \sqrt{\frac{1}{3}}\]

myininaya (myininaya):

now to see if these are inflection numbers we need to test the the concavity around these numbers by choosing test numbers around these numbers and plug them into f''

myininaya (myininaya):

so we know 0 is in between those numbers and we know -5 is before -sqrt(1/3) and 5 is after sqrt(1/3) so we can use these to test the intervals

myininaya (myininaya):

\[f''(-5)=\frac{2((-5)^2+1)(-3(-5)^2+1)}{((-5)^2+1)^4}=\frac{(+)(-)}{(+)}=(-)\] so on the interval (-inf,-sqrt(1/3)) the function is concave down

myininaya (myininaya):

\[f''(0)=\frac{2(+)(+)}{(+)}=(+) \] => (-sqrt(1/3),sqrt(1/3) is concave up

myininaya (myininaya):

i will leave the last interval for you

OpenStudy (anonymous):

the last interval we are using is 5 right?

myininaya (myininaya):

you don't have to use 5 that is just the number i chose after sqrt(1/3)

OpenStudy (anonymous):

okay ill just use 5. so 2((5)^2+1)(-3(5)^2+1)/ ((5)^2+1)^4=(+)(-)/(+) so it is concave downward? right?

myininaya (myininaya):

right :)

OpenStudy (anonymous):

yay! (: do you think you could possible help me with just one more question for calculus?

myininaya (myininaya):

post a new thread and i might get to it okay? :)

OpenStudy (anonymous):

okay thanks! have a great day (: